Badlands, Inc. manufactures a household fan that sells for $25 per unit. All sales are on account, with 40 percent of sales cRequired: 1. Determine the number of units that Badlands sold in December 20x0. 2. Compute the sales revenue for March 20x1.
Answer #1
1 December Sales 8,000 Units
Cash Receipts of December Accounts Receivables in January = $ 120,000
60% of Sales is collected in following Month
December Total sales = $ 120,000 / 60%
= $ 200,000
Selling Price per unit = $ 25
December Sales Unit = 200,000 / 25 = 8,000 units
2 Sales Revenue for March $ 225,000
Expected Sales Unit in March 9,000 units
Selling Price per unit = $ 25
Sales Revenue for March = 9,000 units * $ 25
= $ 225,000
3 First Quarter Sales Revenue $ 630,000
January Sales
Collected in January 85000
Collected in February 135000
220000
February Sales
Collected in February 74000
40% of Sales is collected in same Month
February Sales = $ 74,000 / 40% = $ 185,000
February Sales 185000
March Sales $ 225,000
Total for First Quarter 630000
4 March 31 Accounts Receivable $ 135,000
March 31 Accounts Receivable = March Sales * % to be collected in following month
= 225,000 * 60%
= $ 135,000
5 December 31 Finished Good Inventory 1760 units
Finished Goods Inventory is 20% of next months Sales
January Sales = $ 220,000
Finished Goods Inventory in December = 20% * $ 220,000
= $ 44,000
Selling Price per unit = $ 25
Finished Goods Inventory = $ 44,000 / $ 25
= 1,760 units
6 Finished goods to be manufactured in January 9080 units
Total Sales in January 220000
Total Unit Sales in January ( 220,000 / 25 ) 8800 units
Beginning Inventory 1760 units
10560
Ending Inventory ( 20% * 185000 / $ 25) 1480 units
Finished goods to be manufactured in January 9080 units
7 Financing required in January $ 3,600
Opening Cash 23400
Cash Collected in january 205000
Available Cash 228400
Cash Payments 200000
Closing bal 28400
Minimum Balance to Maintain 32000
Financing Reuired 3600